OpenStudy (anonymous):

When old MacDonald retired he divided up his farm and gave 1/2 the land to his oldest son, 1/3 of the remainder to his daughter and 1/4 of the remainder to his youngest son. He kept the remaining 60 hectares for himself. How much did each child get ? PLEASE SHOW WORK

OpenStudy (anonymous):

OldSonShare = 1/2 = 3/6 DaughterShare = 1/2 * 1/3 = 1/6 YoungSonShare = (1 - (3/6 + 1/6))/4 = (1 - 4/6) / 4 = (2/6) / 4 = 1/12 KidsShares = 6/12 + 2/12 + 1/12 = 9/12 So mcdonalds share is what is left: 3/12 = 1/4 th of the land 1/4th of the land is 60 hectares , so the entire area is 240 hectares. OldSon got 1/2 * 240 hectares = 120 hectares Daugther got 1/6 * 240 hectares = 40 hectares YoungSon got 1/12 * 240 hectares = 20 hectares and mcdonald has 1/4 * 240 hectares = 60 hectares

OpenStudy (anonymous):

THANKS!!! :D